OpenStudy (anonymous):

Math corrections, need help studying im doing ones that are harder than this? can someone solve each of them and explain please

OpenStudy (anonymous):

Write the equation of the line with a slope of -5 and a y-intercept of (0,3). Write the equation of the line with a slope of -1/3 and passing through the point (6, -4). Write the equation of the line passing through the points (0, -4) and (-2, 2). Write the equation of the line passing through the points (-6,1) and (-4,2). Write the equation of the line with an undefined slope, passing through the point (2, 5).

OpenStudy (anonymous):

First thing first, do you know any linear equations?

OpenStudy (anonymous):

For example: y = mx + b Do you have any clue what "m", what "b" stand for?

OpenStudy (anonymous):

yes i do im doing some on my homework at the moment those are ones from a test that i got back b4 spring break i need to practice those

OpenStudy (anonymous):

B would the y intercept if im correct and the m is the points on the graph

OpenStudy (anonymous):

Yes, b ( write strictly in lower case) is y intercept point! But "m" is SLOPE Now can you try the first 2 questions, so I can follow up :)

OpenStudy (anonymous):

y = mx + b called SLOPE - INTERCEPT linear equation form Can you write down other forms here?

OpenStudy (anonymous):

i dnt know how thats why i need someone to solve but explain how to ?

OpenStudy (anonymous):

You know the formula, and in the first question you already had slope m, and b, Can you plug them in?

OpenStudy (anonymous):

y = -5x + 0,3

OpenStudy (anonymous):

b means y coordinate only!

OpenStudy (anonymous):

y = -5x + 3

OpenStudy (anonymous):

Yup :) Now can you proceed to write the FORMULA for POINT-SLOPE form?

OpenStudy (anonymous):

how do i do that

OpenStudy (anonymous):

Look up the formula from your textbook, notes ( You haven't plug any numbers in yet )

OpenStudy (anonymous):

i dnt have any srry notebook is in my locker

OpenStudy (anonymous):

Or... google the formula :)

OpenStudy (anonymous):

what the name of it

OpenStudy (anonymous):

POINT-SLOPE linear form?

OpenStudy (anonymous):

nvr mind is this right ?

OpenStudy (anonymous):

Write the equation of the line with a slope of -1/3 and passing through the point (6, -4) y = mx + b y = -1/3x + -4

OpenStudy (anonymous):

Wrong, because you're using the wrong formula !!!

OpenStudy (anonymous):

hold i got my online txt book

OpenStudy (anonymous):

y - y1 = m(x - x1)

OpenStudy (anonymous):

Yep, see the different between 2 formulas! Now in the second question, slope m and the point coordinate are given

OpenStudy (anonymous):

got it

OpenStudy (anonymous):

Can you plug them in?

OpenStudy (anonymous):

y - (-4) = -1/3[x- (6)]

OpenStudy (anonymous):

Correct, can you simplify the signs?

OpenStudy (anonymous):

y - (-4) = -1/3[x- (6)] y + -4 = -1/3[x- 6] y + -4 = -1/3x- 6

OpenStudy (anonymous):

didnt complete it but am i doing it right

OpenStudy (anonymous):

Always change the sign if in front it is NEGATIVE ! y - ( -4) = y + 4

OpenStudy (anonymous):

y - (-4) = -1/3 (x- 6) y + 4 = -x/3 + 6/3 Do you understand so far, can you continue?

OpenStudy (anonymous):

Don't forget to distribute the -1/3 for ALL the elements inside ( the purpose of parentheses )

OpenStudy (anonymous):

one sec

OpenStudy (anonymous):

like that y - (-4) = -1/3[x- (6)] y + 4 = -1/3x- 6 y + 4 - 4 = -1/3x- 6 - 4

OpenStudy (anonymous):

im going by what my txtbook says

OpenStudy (anonymous):

I don't think you read it correctly!

OpenStudy (anonymous):

Example #1: Write the equation of a line with a slope of 2 and passing through the point (-3,4). We know m = 2 and we have an ordered pair that contains the values x1 = -3 and y1= 4. Plug those values into the point-slope formula, y - y1 = m(x - x1). y - 4 = 2[x - (-3)] Simplify and put the equation of the line in slope-intercept form. y - 4 = 2[x + 3] y - 4 = 2x + 6 y -4 + 4 = 2x + 6 + 4 y = 2x + 10

OpenStudy (anonymous):

Write the equation of a line with a slope of 1/2 and passing through the point (2, -2). We know m = 1/2 and we have an ordered pair that contains the values x1 = 2 and y1= -2. Plug those values into the point-slope formula, y - y1 = m(x - x1) y - (-2) = 1/2 (x - 2) y + 2 = 1/2x - 1 y + 2 - 2 = 1/2x - 1 - 2 y =1/2 x - 3

OpenStudy (anonymous):

..= 2[x + 3] ..= 2x + 6 Did you see how the "2" distribute for both terms inside of the bracket?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

Can you redo this: y + -4 = -1/3[x- 6]

OpenStudy (anonymous):

-1/3[x- (6)] -1/3x- 6

OpenStudy (anonymous):

How about the second term?

OpenStudy (anonymous):

which is that

OpenStudy (anonymous):

subtracting -4 from each side right

OpenStudy (anonymous):

.= 2[x + 3] .. = 2x + 6 In here the second term is "3"

OpenStudy (anonymous):

2

OpenStudy (anonymous):

Redo it, will you: -1/3 ( x - 6) =

OpenStudy (anonymous):

Pay attention to the DISTRIBUTE the NEGATIVE sign, too :(

OpenStudy (anonymous):

-2

OpenStudy (anonymous):

As I just remind you, "Pay attention to the DISTRIBUTE the NEGATIVE sign, too :( "

OpenStudy (anonymous):

i dnt get it

OpenStudy (anonymous):

-1/3 ( x - 6) = - (x/3 - 6/3) = - x/3 + 2

OpenStudy (anonymous):

That's what DISTRIBUTION means!

OpenStudy (anonymous):

or -1/3 ( x - 6 ) = 1/3 ( -x + 6 ) = -x/3 + 2

OpenStudy (anonymous):

can you explain

OpenStudy (anonymous):

Since you just practice with distribution First step: Multiply the NEGATIVE SIGN into both terms Second : Multiply 1/3 for both terms

OpenStudy (anonymous):

ok i see but what do you mean by Multiply the NEGATIVE SIGN into both terms Multiply 1/3 for both terms

OpenStudy (anonymous):

Now you can continue with: y + 4 = = -x/3 + 2

OpenStudy (anonymous):

You don't understand DISTRIBUTION at all, don't you: -1/3 ( x - 6 ) Multiply the NEGATIVE SIGN into both terms: = 1/3 ( -x + 6 )

OpenStudy (anonymous):

ok i see what you mean

OpenStudy (anonymous):

so you move the negative to the x

OpenStudy (anonymous):

AND the second term, too!

OpenStudy (anonymous):

what is the second term

OpenStudy (anonymous):

How many terms inside of the parentheses here???

OpenStudy (anonymous):

1

OpenStudy (anonymous):

If just one term, there's will be no parentheses !!!

OpenStudy (anonymous):

confused what are the terms

OpenStudy (anonymous):

( x - 6 ) first second terms

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so there is x and -6

OpenStudy (anonymous):

-1/3 ( x - 6 ) Multiply the NEGATIVE SIGN into both terms: = 1/3 ( -x + 6 )

OpenStudy (anonymous):

= 1/3 ( -x + -6 )

OpenStudy (anonymous):

ALWAYS change the signs inside of the parenthese, if in front it is NEGATIVE !

OpenStudy (anonymous):

confused

OpenStudy (anonymous):

= 1/3 ( -x + 6 )

OpenStudy (anonymous):

Just opposite!

OpenStudy (anonymous):

like what you did -1/3 ( x - 6 ) Multiply the NEGATIVE SIGN into both terms: = 1/3 ( -x + 6 )

OpenStudy (anonymous):

Yes, that's it!

OpenStudy (anonymous):

ok i see

OpenStudy (anonymous):

so what is my next step

OpenStudy (anonymous):

= 1/3 ( -x + 6 ) Next, Multiply the 1/3 into both terms:

OpenStudy (anonymous):

i though when simplifying you would add the 4 to both sides

OpenStudy (anonymous):

6x 1/3 is 2

OpenStudy (anonymous):

= 1/3 ( -x + 6 ) Next, Multiply the 1/3 into both terms: = -x/3 + 2

OpenStudy (anonymous):

ok i understand that but where did the one go and why is it -x/3

OpenStudy (anonymous):

any number is multiplied by one is itself! In other words, multiply by 1 is doing nothing :/

OpenStudy (anonymous):

2 times 1/3 is 0.66

OpenStudy (anonymous):

2 * 1/3 = 2/3

OpenStudy (anonymous):

2 stays the same!

OpenStudy (anonymous):

i did with a calculator my bad and how does 2 stay the same

OpenStudy (anonymous):

Because multiplied by 1

OpenStudy (anonymous):

wait all you did was multiply 2 by one right

OpenStudy (anonymous):

i get it

OpenStudy (anonymous):

so is that it ?

OpenStudy (anonymous):

= 1/3 ( -x + 6 ) Next, Multiply the 1/3 into both terms: = -x/3 + 2 Do you understand DISTRIBUTION now???